Some of the electrons which are emitted at low velocities from a hot wire (cathode) go through a small hole in a plate that is at a potential of 1000volts with respect to the cathode. What is the velocity of the electrons when they are passing through the hole? A second plate is parallel to the first and 20cm beyond it and is at a potential of -2000volts with respect to the cathode. Describe the motion of electrons in the region between the two plates.


1
Expert's answer
2021-11-17T18:30:45-0500

"qV=mv^2/2\to v=\sqrt{2qV/m}="


"=\sqrt{2\cdot1.6\cdot10^{-19}\cdot1000/(9.1\cdot 10^{-34})}=1.9\cdot10^7\ (m/s)"


In the region between the two plates: "F=ma\to qE=ma\to qV/d=ma\to"


"a=qV/(dm)=1.6\cdot10^{-19}\cdot(-3000)/(0.2\cdot9.1\cdot10^{-31})=-2.6\cdot10^{15}\ (m/s^2)"


In the region between the two plates the electron will move with acceleration


"a=-2.6\cdot10^{15}\ (m/s^2)" .